OpenStudy (anonymous):

How do I divide and simplify 5v-40 divided by v-8/49v

OpenStudy (bahrom7893):

(5v-40)/(v-8)/(49v)

OpenStudy (bahrom7893):

5(v-8)/(v-8)/(49v) = 5/1/(49v) = 5*49V

OpenStudy (bahrom7893):

245v

jhonyy9 (jhonyy9):

5(y-8) ------- = 5*49v = 245v y-8 --- 49v